OpenStudy (anonymous):

5 over 2x + 15 over 2 = 15

OpenStudy (anonymous):

5/2x+15/2=15?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5/2x=15-15/2=15/2 right?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

then 5/x=15 and x=5/15=1/3

OpenStudy (anonymous):

uhh

OpenStudy (anonymous):

idk